In a certain industrial operation 3.54 × 107 g of TiCl4 are reacted with 1.13 × 107 g of Mg.

(a)Calculate the theoretical yield of Ti in grams.

(b)Calculate the percent yield if 7.91 × 106 g of Ti are actually obtained.


1
Expert's answer
2021-03-23T04:22:22-0400

TiCl4 + 2Mg = 2MgCl2 + Ti

TiCl4 is the limiting reagent

TiCl4 = Ti

190g TiCl4 = 48g Ti

3.54 × 107 TiCl4 = (48 × 3.54 ×107 )/ (190)

= 8.94 ×106g

% yield = actual yield/ theoreectical yield ×100%

= 7.91 ×106 / 8.94 ×106 × 100%

= 88.47%